Evaluate each function and simplify where possible.
a. g(-1) if g(x) = -x^2 – x
b. g(x) + 3 if g(x) = x^2 + x
c. h(x – 1) if h(x) = 3x – 2

Answer by Nate(3500)

a. g(-1) if g(x) = -x^2 – x
g(-1) = -(-1)^2 – (-1)
g(-1) = -(1) + 1
g(-1) = 0
b. g(x) + 3 if g(x) = x^2 + x
g(x) + 3 = (x^2 + x) + 3
g(x) + 3 = x^2 + x + 3
c. h(x – 1) if h(x) = 3x – 2
h(x - 1) = 3(x - 1) – 2
h(x - 1) = 3x - 3 – 2
h(x - 1) = 3x - 5